How many grams of water in a cup that requires 4,500J to heat from 20C to 40C?
1
Expert's answer
2020-05-15T12:11:13-0400

Solution:

The energy, Q, needed to increase the temperature of an object of mass m, by a temperature ΔT, where the substance's specific heat is C, is given by:

Q = m × C × ΔT

m = Q / (C × ΔT)


Here,

Q = 4500 J

C = 4.186 J/(gC)

ΔT = Tfinal - Tinitial = 40C − 20C = 20C


Inputting:

m = (4500 J) / (4.186 J g-1∘C-1 × 20C) = 53.75 g

m(H2O) = 53.75 g


Answer: 53.75 grams of water.